OpenStudy (anonymous):

give exact and approximate solutions to three decimal place y^2-14y+49=25

OpenStudy (mimi_x3):

y^2 - 14y +24 = 0 a = 1 , b = 14 , c = -24 Sub it into the quadratic formula.

OpenStudy (mimi_x3):

c=24*

OpenStudy (mertsj):

Factor; (y-2)(y-12)=0 If the product of factors is 0, one of the factors is 0. So y-2 = 0 or y - 12 = 0. So y = 2 or y = 12

OpenStudy (anonymous):

how about this one anyone y^2-6y+9=81

OpenStudy (mertsj):

You really need to learn how to solve these quadratic equations. Factor y^2 - 14y + 49. It is a trinomial square and its factors are (y-7)^2. So now you have (y-7)^2 = 25. Take the square root of both sides and get y - 7 = +/-5. Add 7 to both sides and get 12 or 2.